OpenStudy (anonymous):

3 sqrt 56x ^14 simplify !

OpenStudy (anonymous):

3 sqrt (56x^14) 3 sqrt (2^3 *7 (x^4)^3 *x^2 ) 2x^4 (3 sqrt (7x^2) )

OpenStudy (anonymous):

If I read the question properly

hartnn (hartnn):

i think its \(\huge \sqrt[3]{56x^{14}}\) ??

OpenStudy (anonymous):

Thats what i took it as hartn

OpenStudy (anonymous):

yes it's that hartnn

OpenStudy (anonymous):

Or easier to understand if you write it as ((56)(x^14))^1/3

OpenStudy (anonymous):

How would you simplify it though..?

OpenStudy (anonymous):

well try separate the terms, Leaves you with (56^1/3) and x^14)^1/3.. if you remember the rules of exponents, x^14)^1/3 is simply x^ ( 14*1/3), and 56 is is an irrational number, so you can leave it as (56)^1/3

OpenStudy (anonymous):

oh that's it?

OpenStudy (anonymous):

or, you go further and say 56)^1/3= 2*(7^(1/3)) (since 8 cube root is 2)

OpenStudy (anonymous):

so in the end it would be \[2 \sqrt[3]{7}x^{14/3}\]

OpenStudy (anonymous):

thank you very much